# Mixed-State Measurement-Induced Phase Transitions in Imaginary-Time Dynamics

### v2

@@ III.4 ITE with probabilistic measurements
Now we extend the discussions above to a general case in which each site is independently measured with probability $p$ in each circuit layer.
For a system of $N$ qubits and $k$ circuit layers, the binary choice of measuring or not measuring each qubit in each layer yields $2^{(k-1)N}$ distinct ensembles, which together form an extended ensemble. In this sense, the examples in Figs. 1(b), 2(b), and 3 correspond to three of the $2^{N}$ possible ensembles when $k=2$.
@@ III.5 QMC simulations
Importantly, the diagrammatic representation of the evolving state $\rho_{k}$ naturally facilitates QMC studies of $\rho_{n_{d}}$ based on imaginary-time path integrals (after applying $n_{d}$ layers, we obtain the final output state $\rho_{n_{d}}$), such as the stochastic series expansion (SSE) [Sandvik1992sse, Sandvik1999sse, Sandvik2010lecture, syljuaasen2002quantum, sandvik2019stochastic, Yan2019sweeping, yan2022global], as long as the Hamiltonian is sign-problem-free. In addition, it also makes it straightforward to implement tensor-network–based simulations [Ran2020tensor]. Therefore, we can study the stationary state $\rho_{n_{d}}$ by performing QMC simulations on the generalized partition function $\mathcal{Q}_{n_{d}}\propto\text{Tr}(\rho_{n_{d}})$, which is associated with the extended ensemble.
@@
$\mathcal{Q}_{n_{d}}=\sum_{\{A_{k}\}}\prod_{l=2}^{n_{d}}\bigg[p^{N_{A_{l}}}(1-p)^{N-N_{A_{l}}}\bigg]Q_{n_{d}}(\{A_{k}\})$ (6)
@@
Within the SSE framework, we have developed an unbiased and efficient QMC algorithm to simulate Eq. (6), applicable to Hamiltonians for ITEs in arbitrary spatial dimensions. While the present work focuses on the stationary state $\rho_{n_{d}}$ ($n_{d}\to\infty$), the algorithm also enables studies of the dynamics of the MDITE process. Further details of the QMC algorithm are provided in Appendix A and B.
IV One-dimensional example with discrete symmetry
@@ IV.1 Model and observables
As the first example, we consider the one-dimensional (1D) transverse-field Ising model (TFIM) [pfeuty1970one] as the Hamiltonian in the MDITE protocol, which is given by
$H=-\sum_{i}Z_{i}Z_{i+1}-h\sum_{i}X_{i}$ (7)
@@
This model has a global $\mathbb{Z}_{2}$ symmetry and the ground state exhibits a quantum phase transition at $h_{c,\mathrm{GS}}=1$, as determined by the Kramers-Wannier duality, which belongs to the 2D classical Ising universality class. For $h<h_{c,\mathrm{GS}}$, the system is in a ferromagnetic phase with nonzero magnetization, while for $h>h_{c,\mathrm{GS}}$, it is in a paramagnetic phase with zero magnetization.
Though free energy is not a well-defined observable in the MDITE protocol, the magnetization of spins remains accessible and physical, serving as an important proxy for characterizing the stationary state $\rho_{n_{d}}$. Specifically, for the TFIM, we examine the absolute magnetization $\langle|m|\rangle$, while also considering the second moment $\langle m^{2}\rangle$ and the fourth moment $\langle m^{4}\rangle$, from which the Binder ratio $R_{2}=\langle m^{4}\rangle/\langle m^{2}\rangle^{2}$ is constructed. Note that while the magnetization is defined as $\langle m\rangle:=\langle\sum_{i=1}^{N}Z_{i}\rangle$, $\langle|m|\rangle$ does not correspond to $\langle|\sum_{i=1}^{N}Z_{i}|\rangle$. Instead, in QMC simulations, $\langle|m|\rangle$ represents the ensemble average of the absolute magnetization measured in individual classical QMC samples [Sandvik2010lecture]. For the standard TFIM, observables $\langle|m|\rangle$ and $R_{2}$ are important for identifying the quantum critical point [blote2002cluster, huang2020worm, sandvik2003stochastic, zhou2022quantum, yan2023quantum]. In the ferromagnetic phase, the magnetization distribution is double-peaked around $\pm\langle|m|\rangle$, leading to $R_{2}\to 1$, while in the paramagnetic phase, it becomes Gaussian, yielding $R_{2}\to 3$. Remarkably, as we will show below, these observables can also be used to characterize the mixed-state phase transition of $\rho_{n_{d}}$ in the presence of measurements.
IV.2 Stationary state
@@
First, we show the existence of the stationary state in the MDITE protocol, as noted in Sec. II. Fig. 5 shows that for the control parameters $(\tau,h,p)=(1,1.8,0.66)$—which we will later identify as a critical point—the state $\rho_{n_{d}}$ converges to a stationary state in the limit $n_{d}\to\infty$.
@@
This is evidenced by the rapid convergence of both $\langle m^{2}\rangle$ and the Binder ratio $R_{2}$ with increasing circuit depth. We observe similar convergence behavior for other parameter choices. In practice, we find that a circuit depth of $n_{d}=2L/\tau$ is sufficient to reach the stationary state for the 1D TFIM.
[…]
@@
Figure 5: When setting $(\tau,h,p)=(1,1.8,0.66)$ for the MDITE with the 1D TFIM: (a) Convergence of the second moment $\langle m^{2}\rangle$ with increasing circuit depth for various system sizes $L$. (b) Convergence of the Binder ratio $R_{2}$ with increasing circuit depth for various system sizes $L$.
IV.3 Measurement-induced phase transition
[…]
@@ IV.3 Measurement-induced phase transition
(g) and (h): Critical surfaces viewed from different angles. To facilitate the distinction of critical $p$ corresponding to different parameters on the critical surface, the figure employs a colorbar to map $p$ to different colors.
@@
We explore the three-dimensional parameter space $(\tau,h,p)$ by first fixing $\tau$ and $h$, and then scanning the measurement rate $p$ to determine whether it induces a phase transition. Our extensive numerical calculations reveal that tuning $p$ indeed drives a transition for a wide range of $(\tau,h)$ values when $h>1$.
@@
One example is shown for $(\tau,h)=(1,1.8)$ in Fig. 6(a), where a clear crossing of the Binder ratio $R_{2}$ across different system sizes $L$ identifies a critical point $p_{c}$. The asymptotic behavior of $R_{2}$—tending to 3 for $p<p_{c}$ and to 1 for $p>p_{c}$—provides strong evidence for a mixed-state phase transition analogous to that in the classical Ising model. This indicates a disordered mixed phase below $p_{c}$ and an ordered mixed phase above it.
@@
At $p=0$, the stationary state corresponds to the ground state of the 1D TFIM, which is in the paramagnetic phase for $h=1.8>1$. The results in Fig. 6(a) thus demonstrate that sufficiently strong projective measurements can drive the system into a mixed ferromagnetic phase, polarized along the $z$-axis. In contrast to measurement-induced entanglement-entropy transitions that require quantum trajectory analysis [Skinner2019, buchhold2021effective], this transition is detectable using conventional linear observables without the need for information-theoretic diagnostics.
@@
To identify the universality class of the critical point, we perform a finite-size scaling analysis [privman1990finite] on
$x_{L}=\frac{p-p_{c}}{p_{c}}L^{1/\nu},\quad y_{L}=\langle|m|\rangle L^{\beta/\nu}$ (8)
@@
where $\nu$ and $\beta$ are the critical exponents related to the correlation length and the magnetization. From the data collapse, we extract $p_{c}\approx 0.667$, $\nu\approx 1.08$, $\beta\approx 0.43$, and $\beta/\nu\approx 0.40$, whose stability is reported in Table 1. The corresponding plots are presented in Fig. 6(b). We note that the critical exponents $\nu$ and $\beta$ are quite different from that of the equilibrium 2D or 3D classical Ising criticality, which means this mixed-state phase transition is not a standard Ising phase transition, but within a new universality class.
$L_{\min}$ $p_{c}$ $\nu$ $\beta$ $\beta/\nu$
112 0.6668(6) 1.073(5) 0.422(8) 0.393(5)
128 0.6667(7) 1.077(6) 0.424(9) 0.394(6)
144 0.6665(9) 1.079(7) 0.43(1) 0.395(8)
160 0.666(1) 1.084(8) 0.43(1) 0.397(96)
@@
Table 1: For $(\tau,h)=(1,1.8)$, we perform a data collapse of $\langle|m|\rangle$ using system sizes $L=112,128,144,160,176,192$. To test the stability of the extracted exponents, we first discard the $L=112$ data, setting $L_{\min}=128$, and then progressively increase $L_{\min}$ to $160$ to retain only the three largest system sizes. The resulting stable values are $p_{c}\approx 0.667$, $\nu\approx 1.08$, $\beta\approx 0.43$, and $\beta/\nu\approx 0.40$.
@@ IV.4 $\tau$- and $h$-induced phase transitions
In addition to the measurement-induced transition driven by tuning $p$, the stationary state exhibits mixed-state phase transitions when either the imaginary-time step $\tau$ or the field strength $h$ is varied, with the other parameters fixed. As an example, for $(h,p)=(2.5,0.5)$, the Binder ratio $R_{2}$ shows a clear crossing at $\tau_{c}\approx 0.265$, with a scaling analysis of $\langle|m|\rangle$ yielding $\nu\approx 1.19$ [Fig. 6(c, d), Table 2]. Similarly, for $(\tau,p)=(1.2,0.8)$, we identify a transition at $h_{c}\approx 1.84$, characterized by $\nu\approx 1.18$ and $\beta\approx 0.46$ [Fig. 6(e, f), Table 3].
Although the values of $\nu$ and $\beta$ differ from those of the $p$-driven transitions, the ratio $\beta/\nu\approx 0.40$ remains consistent within error bars across all cases examined. This suggests that $\beta/\nu$ could serve as a universal critical exponent for mixed-state phase transitions in the 1D TFIM under the MDITE protocol with $p\neq 0$. In the following section, we elucidate the mechanism underlying these transitions and demonstrate that they fall within a unique universality class.
$L_{\min}$ $\tau_{c}$ $\nu$ $\beta$ $\beta/\nu$
112 0.2651(1) 1.1873(2) 0.480(1) 0.404(1)
128 0.2652(1) 1.189(1) 0.480(2) 0.404(2)
144 0.2651(2) 1.190(2) 0.480(4) 0.403(3)
160 0.2652(4) 1.191(6) 0.480(7) 0.403(6)
Table 2: For $(h,p)=(2.5,0.5)$, we perform a data collapse of $\langle|m|\rangle$ using system sizes $L=112,128,144,160,176,192$. To test the stability of the extracted exponents, we first discard the $L=112$ data, setting $L_{\min}=128$, and then progressively increase $L_{\min}$ to $160$ to retain only the three largest system sizes. The resulting stable values are $\tau_{c}\approx 0.265$, $\nu\approx 1.19$, $\beta\approx 0.48$, and $\beta/\nu\approx 0.40$.
$L_{\min}$ $h_{c}$ $\nu$ $\beta$ $\beta/\nu$
112 1.8375(3) 1.181(5) 0.46(1) 0.39(1)
128 1.8373(7) 1.18(1) 0.46(1) 0.39(1)
144 1.837(1) 1.19(2) 0.46(1) 0.38(1)
160 1.835(2) 1.21(3) 0.46(1) 0.38(2)
Table 3: For $(\tau,p)=(1.2,0.8)$, we perform a data collapse of $\langle|m|\rangle$ using system sizes $L=112,128,144,160,176,192$. To test the stability of the extracted exponents, we first discard the $L=112$ data, setting $L_{\min}=128$, and then progressively increase $L_{\min}$ to $160$ to retain only the three largest system sizes. The resulting stable values are $h_{c}\approx 1.84$, $\nu\approx 1.18$, $\beta\approx 0.46$, and $\beta/\nu\approx 0.40$.
IV.5 Analysis of the criticalities
The key to understanding these mixed-state phase transitions lies in the dual role of the $Z$-basis: it is both the basis in which projective measurements are performed and the basis in which the ordered phase of the Hamiltonian is defined. Importantly, quantum coherence is generated solely by the off-diagonal terms, namely the transverse field, via the ITE operator $e^{-\frac{\tau}{2}H}$.
In the regime $h>1$ with short imaginary-time steps ($\tau h\ll 1$), frequent projective measurements at high $p$ interrupt the buildup of quantum coherence, producing dynamics reminiscent of the quantum Zeno effect [Misra1977zeno, Itano1990zeno, Koshino2005zeno] when the circuit is sufficiently deep. The transverse field, which drives quantum fluctuations in the $X$-basis, is repeatedly suppressed before correlations can form, as the state is frequently projected onto the $Z$-basis. With off-diagonal coherence effectively eliminated, the dynamics become dominated by the diagonal Ising terms, steering the system toward a ferromagnetic phase analogous to that of the classical Ising model. In contrast, when the transverse field is sufficiently strong or the evolution time is sufficiently long ($\tau h\gg 1$), quantum coherence emerges rapidly within the imaginary-time interval. Coherence is restored faster than it can be suppressed by measurements, preventing the development of ferromagnetic order even at $p=1$. Hence, no phase transition occurs in this regime by tuning the measurement rate $p$.
Between these two limits, the competition between quantum coherence and measurement-induced decoherence establishes a critical balance. Tuning the parameters, including $p$, $\tau$, and $h$, to this balance drives a mixed-state phase transition separating the ordered and disordered phases. In other words, the transitions driven by tuning $p$, $\tau$ or $h$ can be understood through the same mechanism. The distinction lies solely in which parameter modulates the competition between coherent evolution and measurement-induced decoherence.
Building upon the consistent critical exponent ratio $\beta/\nu\approx 0.40$ observed across all dynamical transitions in our simulations, and supported by the unified dynamical mechanism described above, we argue that the mixed-state phase transitions related to the 1D TFIM fall within a single universality class when $p>0$. This universality indicates the existence of a continuous critical surface in the extended parameter space $(\tau,h,p)$ separating the paramagnetic and ferromagnetic phases. According to the finite-size scaling hypothesis and renormalization-group theory [fisher1972scaling, sandvik2010computational, Wilson1971Renormalization, Cardy1996], the magnetization near a critical point obeys the scaling form $\langle m\rangle\sim L^{-\beta/\nu}f(gL^{1/\nu})$, where $g$ is the reduced parameter around the critical point and $f$ is a dimensionless function. In the nonequilibrium universality class considered here, the ratio $\beta/\nu$ remains universal, whereas the individual critical exponents $\beta$ and $\nu$ can depend on the path along which the critical point is approached. Consequently, different parameter trajectories may exhibit distinct values of $\beta$ and $\nu$, while their ratio $\beta/\nu$, which governs the leading finite-size scaling of the magnetization, remains the same [wang2025sudden].
To visualize this structure, we construct a schematic phase diagram based on approximate transition points obtained from Binder ratio crossings in several system sizes, as shown in Fig. 6(g) and (h). Note that this surface serves as a qualitative guide to the phase boundary rather than a precise determination of the critical manifold. Importantly, Fig. 6(h) shows that the critical surface shifts toward larger $\tau$ as $h$ approaches $1$. This trend is expected: longer evolution times require weaker transverse fields to allow ferromagnetic order to emerge under strong measurements. Consequently, in the limit $h\to 1$, the critical surface is expected to extend infinitely along the $\tau$-axis, permitting a transition at any $\tau$ by tuning $p$. A similar reasoning applies in the opposite limit, $h\to\infty$ and $\tau\to 0$.
In addition, the emergence of long-range order with an increasing measurement rate can also be understood qualitatively from the perspective of QMC simulations: a higher measurement rate naturally generates larger clusters that connect a broader range of sites during the QMC updates. This provides a direct and intuitive picture of how measurements drive ordering in the system of TFIM, highlighting the use of the diagrammatic representations in Sec. III and the advantage of QMC in revealing the underlying mechanism. Further details are provided in Appendix. C.
V Two-dimensional example with continuous symmetry
V.1 Model and observables
As the second example, we consider the 2D $S=1/2$ antiferromagnetic columnar dimerized Heisenberg model (CDHM) on a $L\times L$ square lattice [Matsumoto2001qmcdhm]. The Hamiltonian is given by
where $\mathbf{S}_{i}$ is the spin operator at site $i$, and $\langle ij\rangle_{1}$ and $\langle ij\rangle_{g}$ denote two different sets of nearest-neighbor pairs, as shown in Fig. 7. We also consider the local $Z$-basis as the computational basis, and the periodic boundary condition is chosen. The model has a global $SU(2)$ symmetry and the ground state exhibits a quantum phase transition at $g_{c,\mathrm{GS}}\approx 1.90951$ [Sandvik2010lecture, Matsumoto2001qmcdhm, Ma2018Anomalous], which belongs to the 3D $O(3)$ universality class. The critical point separates a Néel ordered phase ($g<g_{c,\mathrm{GS}}$) from a dimerized phase ($g>g_{c,\mathrm{GS}}$).
Due to the antiferromagnetic nature of the Heisenberg interactions, we consider the staggered magnetization $\langle m_{z}\rangle=\Big\langle\sum_{i}(-1)^{x_{i}+y_{i}}Z_{i}\Big\rangle$, where $x_{i}$ and $y_{i}$ denote the integer coordinates of site $i$ in the lattice. As in the TFIM, we compute the absolute magnetization $\langle|m_{z}|\rangle$, the second and fourth moments of the magnetization, and the Binder ratio $R_{2,z}$. Furthermore, for the ground state of the CDHM, the Binder ratio approaches $R_{2,z}\to 9/5$ in the Néel phase and $R_{2,z}\to 3$ in the dimerized phase [sandvik1997finite, wang2006high, ding2018engineering, Ma2018Anomalous].
V.2 Stationary state
As in the 1D TFIM case, the MDITE protocol in this case also drives the system toward a stationary state in the limit of large circuit depth. Fig. 8 illustrates this convergence for representative parameters $(\tau,g,p)=(1,3.5,0.5)$, where both the staggered magnetization $\langle|m_{z}|\rangle$ and the Binder ratio $R_{2,z}$ quickly approach their stationary values. In practice, we again find that a circuit depth of $n_{d}=2L/\tau$ is sufficient to reach the stationary state, and this choice is used in all subsequent QMC simulations for the CDHM.

### v3

@@ III.4 ITE with probabilistic measurements
Now we extend the discussions above to a general case in which each site is independently measured with probability $p$ in each time step.
For a system of $N$ qubits and $k$ time steps, the binary choice of measuring or not measuring each qubit in each layer yields $2^{(k-1)N}$ distinct ensembles, which together form an extended ensemble. In this sense, the examples in Figs. 1(b), 2(b), and 3 correspond to three of the $2^{N}$ possible ensembles when $k=2$.
@@ III.5 QMC simulations
Importantly, the diagrammatic representation of the evolving state $\rho_{k}$ naturally facilitates QMC studies of $\rho_{n_{d}}$ based on imaginary-time path integrals (after applying $n_{d}$ layers, we obtain the final output state $\rho_{n_{d}}$), such as the stochastic series expansion (SSE) [77, 79, 81, 93, 82, 116, 118], as long as the Hamiltonian is sign-problem-free. In addition, it also makes it straightforward to implement tensor-network–based simulations [75].
@@
$\mathcal{Q}_{n_{d}}=\sum_{\{A_{k}\}}\prod_{l=2}^{n_{d}}\bigg[p^{N_{A_{l}}}(1-p)^{N-N_{A_{l}}}\bigg]Q_{n_{d}}(\{A_{k}\}),$ (9)
@@
Within the SSE framework, we have developed an unbiased and efficient QMC algorithm to simulate Eq. (9), applicable to Hamiltonians for ITEs in arbitrary spatial dimensions. While the present work focuses on the steady state $\rho_{\infty}$ ($n_{d}\to\infty$), the algorithm also enables studies of the dynamics of the MDITE process. Further details of the QMC algorithm are provided in Appendices A and B.
IV Numerical results for a 1D Model with discrete symmetry
@@ IV.1 Model and observables
As the first example, we consider the one-dimensional (1D) transverse-field Ising model (TFIM) [72] as the Hamiltonian in the MDITE protocol, which is given by
$H=-\sum_{i}Z_{i}Z_{i+1}-h\sum_{i}X_{i},$ (10)
@@
This model has a global $\mathbb{Z}_{2}$ symmetry and the ground state exhibits a quantum phase transition at $h_{c,\mathrm{GS}}=1$, as determined by the Kramers-Wannier duality, which belongs to the 2D classical Ising universality class. For $h<h_{c,\mathrm{GS}}$, the system is in a ferromagnetic (FM) ordered phase with nonzero magnetization, while for $h>h_{c,\mathrm{GS}}$, it is in a paramagnetic (PM) disordered phase with zero magnetization.
Though free energy is not a well-defined observable in the MDITE protocol, the magnetization of spins, $\langle m\rangle\equiv\langle\sum_{i=1}^{N}Z_{i}/N\rangle$, remains accessible and physical, serving as an important diagnosis for characterizing the steady state $\rho_{\infty}$.
Specifically, for the TFIM, we numerically calculate the absolute magnetization $\langle|m|\rangle$, while also considering the second moment $\langle m^{2}\rangle$ and the fourth moment $\langle m^{4}\rangle$, from which the dimensionless Binder ratio $R_{2}=\langle m^{4}\rangle/\langle m^{2}\rangle^{2}$ is obtained. Note that $\langle|m|\rangle$ does not correspond to $\langle|\sum_{i=1}^{N}Z_{i}/N|\rangle$. Instead, in QMC simulations, $\langle|m|\rangle$ represents the ensemble average of the absolute magnetization measured in individual classical QMC samples [81]. For the standard TFIM, observables $\langle|m|\rangle$ and $R_{2}$ are important for identifying the phase transition point from finite-size simulations [4, 36, 80, 128, 117].
In the FM phase, the magnetization distribution is double-peaked around $\pm\langle|m|\rangle$, serving as a direct signature of spontaneous symmetry breaking (SSB). This behavior leads to $R_{2}\to 1$, while in the PM phase the distribution becomes Gaussian, yielding $R_{2}\to 3$. Remarkably, as we will show below, these observables can also be used to characterize the mixed-state phase transitions of $\rho_{\infty}$ in the presence of measurements.
The current section mainly presents the numerical results obtained from QMC simulations. The corresponding discussions and theoretical analysis will be provided in Sec. V.
IV.2 Steady state
@@
We here demonstrate the existence of the steady state in the MDITE protocol, as noted in Sec. II. Fig. 5 shows that for the control parameters $(\tau,h,p)=(1,1.8,0.66)$—which we will later identify as a critical point—the state $\rho_{n_{d}}$ converges to a steady state in the limit $n_{d}\to\infty$.
@@
This is evidenced by the rapid convergence of both $\langle m^{2}\rangle$ and the Binder ratio $R_{2}$ with increasing the number of time steps, producing dynamics reminiscent of the quantum Zeno effect [62, 39, 42]. We observe similar convergence behavior for other parameter choices.
In practice, we find that $n_{d}=2L/\tau$ is sufficient to reach the steady state for the 1D TFIM, which implies that the dynamical exponents for the mixed-state critical points satisfy that $z\leq 1$. As we will show in Sec. IV.5, our numerical simulations give $z\approx 1$ for all the critical points investigated here.
@@
Figure 5: When setting $(\tau,h,p)=(1,1.8,0.66)$ for the MDITE with the 1D TFIM: (a) Convergence of the second moment $\langle m^{2}\rangle$ with increasing the number of time steps for various system sizes $L$. (b) Convergence of the Binder ratio $R_{2}$ with increasing time steps for various system sizes $L$.
IV.3 Mixed-state MIPT
@@
We first explore the three-dimensional parameter space $(\tau,h,p)$ by fixing $\tau$ and $h$, and then scanning the measurement rate $p$ to examine whether it induces a phase transition. Our extensive numerical calculations show that tuning $p$ indeed drives a transition for a wide range of $(\tau,h)$ values when $h>1$.
@@
One example is shown for $(\tau,h)=(1,1.8)$ in Fig. 6(a), where a clear crossing of the Binder ratio $R_{2}$ for different system sizes $L$ identifies a critical point $p_{c}$. The asymptotic behavior of $R_{2}$—approaching $3$ for $p<p_{c}$ and $1$ for $p>p_{c}$—provides strong evidence for a mixed-state phase transition. Moreover, since the measurement channel preserves the weak (or average) $\mathbb{Z}_{2}$ symmetry of the state [76], the observed transition supported by the behavior of binder ratios thus corresponds to a weak-to-nothing SSB, analogous to thermal phase transitions in the 2D classical Ising model. This separates a disordered mixed phase for $p<p_{c}$ and an ordered mixed phase for $p>p_{c}$.
@@
Note that when $p=0$, the steady state reduces to the ground state of the 1D TFIM, which is in the PM phase when $h=1.8>1$. The results in Fig. 6(a) thus demonstrate that sufficiently strong projective measurements can drive the system into a mixed FM phase, polarized along the $z$-axis. In contrast to measurement-induced entanglement-entropy transitions that require quantum trajectory analysis [89, 6], this transition is detectable using conventional linear observables without the need for information-theoretic diagnostics.
@@
To identify the universality class of the critical point, we perform a finite-size scaling analysis [74] on
$x_{L}=\frac{p-p_{c}}{p_{c}}L^{1/\nu},\quad y_{L}=\langle|m|\rangle L^{\beta/\nu},$ (11)
@@
where $\nu$ and $\beta$ denote the critical exponents related to the correlation length and the magnetization. From the data collapse, we extract $p_{c}\approx 0.667$, $\nu\approx 1.08$, and $\beta\approx 0.43$, whose numerical stability is reported in Table 1. The corresponding plots are presented in Fig. 6(b).
Importantly, the critical exponents $\nu$ and $\beta$ are quite different from that of the equilibrium 2D or 3D classical Ising criticality, which means this mixed-state phase transition is not a standard Ising phase transition, though the numerical results indicate an Ising-like $\mathbb{Z}_{2}$ SSB.
Parameters $L_{\min}$ Critical point $\nu$ $\beta$
$(\tau,h)=(1,1.8)$ 112 $p_{c}=0.6668(6)$ 1.073(5) 0.422(8)
128 $p_{c}=0.6667(7)$ 1.077(6) 0.424(9)
144 $p_{c}=0.6665(9)$ 1.079(7) 0.43(1)
160 $p_{c}=0.666(1)$ 1.084(8) 0.43(1)
$(h,p)=(2.5,0.5)$ 112 $\tau_{c}=0.2651(1)$ 1.1873(2) 0.480(1)
128 $\tau_{c}=0.2652(1)$ 1.189(1) 0.480(2)
144 $\tau_{c}=0.2651(2)$ 1.190(2) 0.480(4)
160 $\tau_{c}=0.2652(4)$ 1.191(6) 0.480(7)
$(\tau,p)=(1.2,0.8)$ 112 $h_{c}=1.8375(3)$ 1.181(5) 0.46(1)
128 $h_{c}=1.8373(7)$ 1.18(1) 0.46(1)
144 $h_{c}=1.837(1)$ 1.19(2) 0.46(1)
160 $h_{c}=1.835(2)$ 1.21(3) 0.46(1)
@@
Table 1: Critical exponents obtained from data collapse of $\langle|m|\rangle$ with $L=112,128,144,160,176,192$. To test the stability of the extracted exponents, we first discard the $L=112$ data, setting $L_{\min}=128$, and then progressively increase $L_{\min}$ to 160160 to retain only the three largest system sizes.
@@ IV.4 $\tau$- and $h$-induced phase transitions
In addition to the MIPTs driven by tuning $p$, the steady state exhibits mixed-state phase transitions when either the imaginary-time step $\tau$ or the field strength $h$ is varied, with the other parameters fixed.
As an example, for $(h,p)=(2.5,0.5)$, the Binder ratio $R_{2}$ shows a clear crossing at $\tau_{c}\approx 0.265$, with a scaling analysis of $\langle|m|\rangle$ yielding $\nu\approx 1.19$, $\beta\approx 0.48$ [Fig. 6(c, d), Table 1].
Similarly, for $(\tau,p)=(1.2,0.8)$, we identify a transition at $h_{c}\approx 1.84$, characterized by $\nu\approx 1.18$ and $\beta\approx 0.46$ [Fig. 6(e, f), Table 1]. Notably, the values of $\nu$ and $\beta$ change for different critical points.
IV.5 Dynamical scaling
To better understand these mixed-state phase transitions, we also extract the dynamical exponent $z$ by simulating the relaxation dynamics from the initial state to the steady state. Specifically, we determine the relaxation time $\xi_{\tau}$ by fitting the convergence of the order parameter $\langle|m_{z}|\rangle$ as a function of the number of time steps $n_{d}$ using the fitting formula
$m(n_{d})=a\exp(-n_{d}/\xi_{\tau})+b,$ (12)
for each system size. At the critical point, the relaxation time is expected to scale as $\xi_{\tau}\propto L^{z}$.
For the measurement-induced critical point $(\tau,h,p)=(1,1.8,0.6667)$ in Sec. IV.3, we obtain $z\approx 1$. The data are shown in Fig. 7, and the stability analysis yields consistent results, as summarized in Table 2.
Figure 7: When setting $(\tau,h,p)=(1,1.8,0.6667)$ for the MDITE with the 1D TFIM: (a) The relaxation process of $\langle|m|\rangle$ with increasing the number of time steps $n_{d}$ for various system sizes $L$. It satisfies the formula $m=a\exp(-n_{d}/\xi_{\tau})+b$, where $a$, $\xi_{\tau}$ and $b$ are fitting parameters. (b) Extracting the dynamical exponent $z$ from $\xi_{\tau}$ with various system size $L$. The fitting formula is $\xi^{-1}_{\tau}=kL^{-z}$, where $k$ and $z$ are fitting parameters.
$L_{\min}$ $k$ $z$
64 7.1(2) 0.997(5)
72 6.9(2) 0.993(6)
80 6.7(3) 0.987(7)
88 6.5(3) 0.978(9)
96 6.8(4) 0.99(1)
Table 2: For $(\tau,h,p)=(1,1.8,0.6667)$, we extract the dynamical exponent $z$ from the relaxation correlation length using system sizes L=64,72,,128L=64,72,...,128. To test the stability of the extracted exponent, small system sizes are gradually discarded. $L_{\min}$ denotes the smallest systems size used in each fitting.
Similarly, we extract the dynamical exponents for the other two parameter sets $(\tau,h,p)=(1.2,1.8373,0.8)$ and $(0.4675,2.5,0.5)$ in Sec. IV.4, where the phase transitions are driven by the transverse field $h$ and the imaginary-time step $\tau$, respectively. As reported in Appendix F, the exponents are also close to one.
After accounting for finite-size effects, our numerical results are consistent with a dynamical critical exponent $z\approx$1, suggesting an emergent space–time isotropy at the steady-state critical points. This behavior is compatible with an emergent Lorentz-invariant description and possibly an effective conformal field theory, although further evidence is required to substantiate this interpretation.
V Analysis of the criticalities in MDITE steady states
Let us now analyze the novel mixed-state phase transitions observed in the numerical simulations presented in the previous section.
Note that there are three tractable continuous limits: the continuous-time limit ($\tau\to 0$, $p\sim O(1)$), the weak-measurement limit ($p\to 0$, $\tau\sim O(1)$), and the combined limit ($\tau\to 0$, $p\to 0$, with $\tau\sim p$). Away from these limits, analytical description could be challenging.
As we show below, all three limits give rise only to the conventional Ising universality class. Particularly, neither the continuous-time limit nor the weak-measurement limit yields results that are even qualitatively consistent with the numerical findings obtained away from these regimes. For this reason, we defer detailed perturbative calculations and discussions of these two limits to Appendices C and D, respectively.
By contrast, the combined limit provides qualitatively consistent behavior, although the resulting universality class remains that of the standard Ising transition. We thus conclude that, sufficiently far from these continuous limits (such as the numerical results in Sec. IV), the mixed-state phase transitions may fall outside the scope of known universality classes.
V.1 Choi representation
We first represent the steady state in the double-Hilbert space formalism. Using the identities
[…]
the local measurement channel in Eq. (3) can be rewritten as
$\mathcal{E}_{p,i}(\rho)=\bigg(1-\frac{p}{2}\bigg)\rho+\frac{p}{2}Z_{i}\rho Z_{i}.$ (14)
Under the Choi-Jamiołkowski isomorphism [13, 40], $\mathcal{E}_{p,i}$ can be mapped onto a superoperator acting on a doubled Hilbert space, which is
$\mathcal{E}_{p,i}\longrightarrow\hat{\mathcal{E}}_{p,i}=(1-p)^{1/2}e^{\tilde{\tau}\,Z_{i}^{L}Z_{i}^{R}},$ (15)
where the inter-copy coupling strength is given by
$\tilde{\tau}\equiv\tilde{\tau}(p)=\tanh^{-1}\!\left(\frac{p}{2-p}\right),$ (16)
and superscripts $L$ and $R$ denote operators acting on the left and right copies of the Hilbert space, respectively. In addition, for $x\equiv p/(2-p)\in[0,1]$, we can rewrite
$\tilde{\tau}=\tanh^{-1}(x)=\frac{1}{2}\ln\bigg(\frac{1+x}{1-x}\bigg)=-\frac{1}{2}\ln(1-p).$ (17)
When $\tilde{\tau}\to 0$, $\tilde{\tau}$ can be approximately replaced by $p/2$.
Accordingly, the evolution described in Sec. II becomes
$\left|\rho_{k}\right\rangle\!\rangle\propto e^{-\tau\tilde{H}}e^{\tilde{\tau}\sum_{i}Z_{i}^{L}Z_{i}^{R}}\left|\rho_{k-1}\right\rangle\!\rangle,$ (18)
with the two-copy Hamiltonian defined as
$\tilde{H}\equiv\frac{1}{2}(H^{L}+H^{R}).$ (19)
V.2 Effective Hamiltonian in the combined limit
In the limit $\tau\to 0$ and $\tilde{\tau}\to 0$, with $\tau\sim\tilde{\tau}$, Eq. (18) can be simplified using the Baker-Campbell-Hausdorff (BCH) formula. Keeping terms up to linear order in $\tau$ and neglecting $O(\tau^{2})$ corrections, we obtain
$e^{-\tau\tilde{H}}e^{\tilde{\tau}\sum_{i}Z_{i}^{L}Z_{i}^{R}}\approx e^{-\tau H_{\text{eff}}},$ (20)
where the effective Hamiltonian is given by
$H_{\text{eff}}\equiv\tilde{H}-\frac{p}{2\tau}\sum_{i}Z_{i}^{L}Z_{i}^{R}.$ (21)
Therefore, the ITE and measurement operations can be combined into a single effective transformation in this limit, which is exactly an doubled-space ITE with time length $\tau$, under an effective two-copy Hamiltonian $H_{\mathrm{eff}}$.
Moreover, the initial state in the doubled space is given by